*Again-Rules for rounding: If you are calculating in tablets, always round to the nearest 1/2 tablet....

  1. *Again-Rules for rounding:
    • If you are calculating in tablets, always round to the nearest 1/2 tablet.
    • When you are calculating IV' rates, round to the nearest whole number.
    • Perform rounding only the end of the calculation
  2. CONVERSTIONS TO USE
  3. 1 grain =60 mg
  4. 15 grain=1 gram
  5. 1 pint= 2 cups=480mL
  6. 1 quart= 2 pt= 4 cups=960mL
  7. 1 oz= 30 mL
  8. 1 cup = 8 oz = 240 mL
  9. 1 quart = 2 pt = 4cups= 960 mL
  10. 1cc= 1mL
  11. 1tsp= 5mL
  12. 1 Tbs= 15 mL
  13. 1 gallon= 4 qt = 8 pt
  14. Note:Some drug conversions will use cc(cubic centimeters) This is the same as a mL. 1cc=1mL

(*There are actually 29.57 mL in 1 fl oz, but 30 mL is usually used by nurses in medical facilities. We will stay with this standard for all liquid measurements. 1 oz=30mL, 1 c=240mL, 1 pt=480mL, 1 quart=960mL.)

QUESTIONS:

1) Mr. Hill has the following prescription: Ativan (lorazepam) 4 mg PO daily; divide the total dose so that equal amounts are given every 6 hours. You mark the times on the medication sheet according to hospital policy to be 0600, 1200, 1800, and 2400 hours. What is the dose of Ativan to be given at each of those times?

Group of answer choices

4mg

1mg

2mg

6.6mg

2) A client admitted to the hospital with acute renal failure typically takes moexipril (Univasc) for high blood pressure. The client’s BUN is elevated and the pharmacy recommends a 50% dose reduction for this medication. The client’s typical dose is 7.5 mg. The nurse would expect the physician to order:

Group of answer choices

5.75 mg

3.75 mg

2.25 mg

4.25 mg

3) While working in the pediatric clinic, the nurse receives a call from a client trying to figure out how much infant acetaminophen (Tylenol) to give to her child. The doctor has prescribed 45 mg every 4 hours. The medication is provided in 60 mg/5 mL concentration. How much medication (in mL) should the client provide per dose? (Round to the nearest mL)________mL
Answer #1

1.every 6 hrs in 24 hours means 4 times in one day .Mr.hill have a prescribed dose of 4 mg P/O daily ,so we have to divide the total dose by 4 so it will be 4/4=1 mg ,

06.00 =1 mg

12.00=1 mg

18.00=1 mg

24.00=1 mg

By this process of timing Mr.hill can get total. 4 mg of tab.ativan in a day.

2. The client total dose is 7.5 mg

So 50% dose reduction means half of the total dose that is 7.5/2=3.75mg

So the nurse would expect the physician to order 3.75 mg.

3.The medication is provided 60mg/5 ml

5 ml of medication contain 60 mg

So 1 ml medication =60/5=12 mg

As per doctors prescription 45 mg of acetaminophen(Tylenol)

have to give to the child every 4 hours.

So we can calculate the dose in ml by this formula=

12 mg medicine present in 1 ml

1mg medicine present in 1/12 ml

45 mg medicine present in 1/12*45ml

The answer is=1/12*45=3.749~3.75(by approximation method in math)~4 ml (round to nearest)